### v1

[…]
These authors contributed equally to this work.
[…]
These authors contributed equally to this work.
[…]
Characterizing the Reasoning Abilities of Large Language Models via 3-SAT Phase Transitions// The Ability to Reason: A Characterization of LLMs via 3-SAT
[…]
keywords:
keyword1, Keyword2, Keyword3, Keyword4
1 Introduction
The success and versatility of Large Language Models (LLMs) has sparked a widespread interest in their reasoning abilities. (Luc) Let’s mix that sentence There are different views on what reasoning in this context means, whether it is oriented towards common-sense reasoning (REFS) or towards logical or deductive reasoning. We will adhere to Leon Bottou’s definition, which states that reasoning is “[the] algebraic manipulation of previously acquired knowledge in order to answer a new question”. Recent studies suggest that LLMs are inherently capable of zero-shot reasoning [llms_zero_shot_reasoners], that is, performing multi-step inference processes in previously unseen situations. Combined with smart prompting mechanisms that encourage LLMs to think step-by-step, the reasoning abilities of LLMs can be enhanced [chain_of_thought, leasttomost, react, adapt]. Demonstrations of the reasoning abilities of LLMs include but are not limited to planning [llms_zero_shot_planners, saycan, progprompt, codeaspolicies, innermonologue, saycanpay], theorem proving [jiang2023draft, natural_prover, verifier_guided_proof], search and optimization [llms_as_optimizers, llmatic, fun_search], self-reflection [tree_of_thoughts, selfrefine, reflexion], API calling [hugginggpt, toolformer], etc.
On the other hand, there is a significant body of research that presents a critical view of the reasoning abilities of LLMs. These studies highlight cases where LLMs struggle to provide the correct answers. For instance, LLMs may exhibit limitations in consistent logical reasoning [gpt4_cant_reason, llms_are_greedy_reasoners], effective planning [llms_cant_plan], and conducting accurate self-assessment of their outputs [gpt_graph_coloring]. These contrasting perspectives create an ambiguous picture about the true reasoning abilities of LLMs.
According to Leon Bottou [13] reasoning is “[the] algebraic manipulation of previously acquired knowledge in order to answer a new question”.
2 3-SAT problems
To clarify this picture, we evaluate the performance of LLMs on solving 3-SAT problems. 3-SAT is one of the most fundamental problems in computer science and reasoning. The 3-SAT problem is introduced via the SAT-game in BOX 1 (and state them also in implicative form).
It is the prototypical NP-complete problem lying at the basis of computational complexity. To date it is unknown whether there exist efficient (that is, polynomial time) algorithms for NP-complete problems. Numerous important computational problems are NP-complete and can be efficiently reduced to 3-SAT. Furthermore, any polynomial time algorithm for 3-SAT would result in an efficient algorithm for these other NP-complete problems. One other such NP-complete problem is $k$-SAT, which generalizes 3-SAT by lifting the restriction of 3 literals per clause. SAT-solvers are frequently used to solve reasoning problems. First, $k$-SAT problems are constraint satisfaction problems (over boolean variables). Second, SAT-solvers are used to prove that a sentence $s$ is logically entailed by a set of clauses $S$ (notation $S\models s$) using the principle of reductio ad absurdum, which corresponds to adding the negation of the rule $NOT~{}s$ to the set of clauses $S$ and showing that $S~{}AND~{}(NOT~{}s)$ is unsatisfiable. Third, various other reasoning problems in planning and constraint satisfaction can be solved by reducing them to SAT-solvers.
$k$-SAT is well-studied, both theoretically and empirically. The later is especially relevant to obtain insight into the ability of LLMs to reason. It is well-known that there are easy and hard problem instances in 3-SAT (SELMAN et al). Considering the 3-SAT game with a fixed number of variables in Box 1, the number of solutions decreases with the number of clauses. Figure 1 depicts the probability that there exists a satisfying assignment for a particular problem as a function of $N/V$ where $V$ is the number of variables and $N$ the number of rules. It also shows the the phase transition that divides the space into the typically easy - hard - easy regions. When there are only few rules, it is very likely that there exists a solution and it is easy to find it. When there are very many rules, it is unlikely that a solution exists, and close to the phase transition one finds the really hard problems. By analogy to the use of these insights to assess the difficulty of SAT and the performance of solvers, we will use them to empirically assess the reasoning abilities of LLMs.
(Luc) There are also the theoretical arguments (about complexity) that we should write a sentence about. (Gabriele) I’m not entirely sure what theoretical arguments you are referring to..? If you clarify I can try to write something about it. Also, I think something is missing from the story to convince the readers that 3-SAT is a good candidate to do this reasoning-capability study for LLMs. Maybe I would stress a bit more the importance of the NP class. Possibly, you were referring to this?
3 Evaluating the reasoning ability of LLMs on 3-SAT
To evaluate the extent to which LLMs can solve 3-SAT problems, we consider both the decision and the search variant of the 3-SAT problem. In the 3-SAT decision problem, the solver is given a 3-SAT formula and has to return YES when there exists a satisfying assignment and NO when there is no such assignment. In the search version of 3-SAT, the solver must return a satisfying assignment if such an assignment exists, and NO otherwise.
We are seeking answers to the following experimental questions:
(1) Which type of 3-SAT problems can LLMs solve?
(2) Can LLMs use a solver to enhance their reasoning ability?
(3) Are there differences in the solving ability of state-of-the-art LLMs?
We use GPT-4 as the reference LLM, given its superior performance in existing benchmarks [^1] https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard. Subsequently, to answer question (3) we extend our analysis to include comparisons with various other state-of-the-art LLMs.
Furthermore, as the type of input and the prompting strongly influence the results obtained with LLMs, we consider two types of input: (i) SAT-Menu uses a menu-selection scenario to which the CNF formula has been mapped (see Figure X), and (ii) SAT-CNF uses the CNF formula directly as input (see Appendix Table C), and we explore various prompting techniques incorporating in-context learning [llms_few_shot_learners] and chain-of-thought (CoT) [chain_of_thought] prompts. Moreover, we always instruct the LLMs to think step-by-step [llms_zero_shot_reasoners].
Which type of 3-SAT problems can LLMs solve?
  • 3-SAT Decision Problem: As shown in the confusion matrix (Figure 3 [Right]), GPT-4 shows proficiency in identifying unSAT scenarios, though it occasionally misclassifies SAT problems as unSAT, particularly within the Hard region of $\alpha$.
  • 3-SAT Search Problem: Investigating GPT-4’s handling of the 3-SAT search problem reveals a phase transition pattern similar to that of conventional solvers, characterized by Easy-Hard-Easy phases (Figure 2). While GPT-4 demonstrates competent performance in the Easy region, its accuracy significantly diminishes to $\approx 10\%$ in the Hard region. This pattern underscores the model’s variable effectiveness across the phase transition – how GPT-4 is significantly better at reasoning through inherently easy problems than with hard problems.
  • Size of solution space: In Figure 3[Left], we also plot the performance against the satisfiability ratio, given as model $\frac{\text{model count}}{2^{n}}$. Here, model count gives the number of feasible solutions and $n$ is the number of variables. Thus the satisfiability ratio denotes the probability that a randomly selected variable assignment satisfies the given 3-SAT theory. The higher this probability the more likely it is that GPT-4 finds a solution.
  • Impact of Input Type on Reasoning: We observed no change in performance with direct SAT-CNF inputs over SAT-Menu (Figure 9). Thus, the phase transitions remain the same regardless of the input.
  • Impact of Prompt Engineering: Prior works have shown that in-context learning can enhance LLM performance [llms_few_shot_learners]. We explored this with 3-shot learning using three input/output (I/O) examples in the prompt. These examples were randomly selected from a set of accurate solutions generated by the LLM and manually checked for consistency between solutions and their explanations. Each I/O example included both the input preferences and the output solution along with its explanation. From Figure 4 [Left], we saw performance gains in the initial Easy-Hard phase but observed a decrease in the subsequent Easy phase. (Rishi) Explain …
(Luc) What is our explanation / conclusion for these experiments ? Our findings reveal that irrespective of prompt engineering techniques, GPT-4 struggles in reasoning, especially in the Hard region. Even within the Easy region, GPT-4’s performance drops if the satisfiability ratio is low.
(Luc) The differences between the decision problem and solving the SAT search problem is not always clear in the graphs.
Can LLMs use a solver to enhance their reasoning ability?
Here we feed the LLM with the SAT-Menu input and prompt it to map it to a CNF formula. The resulting approach, termed SAT-Translate, thus uses an LLM to translate a natural language task into a solver-compliant SAT formula and then applies a SAT solver for deriving a solution (see Appendix Box D). This underscores the synergy between the LLMs’ capabilities and the solvers’ computational precision [word_to_world_models, dinu2024symbolicai, llmp, linc, xie2023translating, thor_theorem_proving, mathematical_conjecturing_with_llms]. As shown in Figure 4 [Right], the performance is almost perfect. This success is attributed to the relatively lower complexity of translation tasks compared to the NP-complete nature of solving SAT-problems.
Are there differences in the solving ability of state-of-the-art LLMs?
We have conducted comparative experiments with some of the best and most advanced LLMs. This includes open-source (Llama 2 70B (chat-hf) [meta_llama2], Mixtral [mistral_mixtral]) as well as proprietary models (GPT-4 (Turbo) [openai_gpt4], GPT-3.5 (Turbo) [openai_gpt3_5], Gemini Pro [google_gemini], PaLM 2 (text-bison) [google_palm]). In these evaluations, GPT-4 emerges as a notable exception, outshining other LLMs across various metrics (cf. the accompanying material).
We identify both common and divergent performance patterns.
Common Performance Trends:
  • Size of solution space: Across LLMs, performance correlates with solution space size (Figure 8).
  • SAT-Menu vs. SAT-CNF: In general, LLM performance does not change with input type (Figure 9).
  • SAT-Translate: Integration of LLMs with solvers is more effective (Figure 9).
Divergent Performance Trends:
  • Phase Transitions: As shown in Appendix Figure 7 [Left], GPT-4 uniquely reflects solver-like phase transitions. In contrast, other LLMs exhibit a marked decline in performance in the high $\alpha$ region, displaying an Easy-Hard-Hard pattern. This divergence primarily stems from their failure to identify unSAT problems, which predominate at higher $\alpha$ values (see Appendix Figure 7 [Right]).
  • In-context improvements: While GPT-4 shows improvements, other LLMs show no performance change with in-context learning (Figure 10).
3.1 Can LLMs use a solver to enhance their reasoning ability?
BOX 2: show the inputs for a LLM.
Figure 1: The SAT Game is a gamification of the SAT problem. In the SAT game, one must find a valuation of Boolean variables (true/false) such that a Boolean CNF formula evaluates to true. A line in the game actually represents a clause, which is a disjunction of literals (variables or their negation connected by logical OR). A clause evaluates to true as soon as one of its literals is true. Since clauses are conjunctively connected (by a logical AND), all clauses must evaluate to true so that the formula be true.
4 Related Works
We group the papers into three categories – Papers that perform an in-depth exploration of reasoning abilities; papers that perform an empirical evaluation of reasoning abilities on benchmark datasets and tasks; and finally papers that propose methods to improve reasoning.
Intensive Studies
Recent works have examined how transformer architecture, training, and inference paradigms influence the reasoning abilities of LMs. Training analysis shows how LMs like Bert can fit on statistical features, instead of emulating correct reasoning, to achieve high performance [paradox_of_learning_to_reason]. This is also linked to the Clever Hans Cheat in teacher-forcing training where LLMs identify shortcuts that bypass true reasoning [pitfalls_of_next_token_prediction]. In terms of inference, the autoregressive nature of LLMs make them prone to snowballing errors over time [dziri2023faith]. A more systematic approach is to directly evaluate the predicted chains of thought, wherein it has been revealed how LLMs are prone to generating unfaithful and biased explanations [llms_dont_say_what_they_think]. Moreover, LLMs reason greedily, failing in contexts with multiple valid reasoning steps [llms_are_greedy_reasoners].
On the architectural side, recent research has shown that the transformer layer is incapable of function composition for larger domains [limits_of_transformer_architectures, dziri2023faith]. Moreover, it has been shown that the computation of transformers lies in the complexity class log-uniform TC0, which is more restrictive than the logarithmic memory class L [log_precision_transformers]. Given that the $2-$SAT problem is NL-complete (i.e., the class of non-deterministic logarithmic space algorithms), multi-layer transformers cannot solve $2-$SAT instances, unless L$=$NL [limits_of_transformer_architectures].
In this work, we delve into reasoning [^2] We specifically focus on search-based reasoning while excluding domains like commonsense reasoning, an emergent ability of LLMs, shown to manifest with scale [wei2022emergent, llms_few_shot_learners, bigbench]. However, existing works have also brought this claim to question, arguing if the emergent abilities are but a mirage stemming from choice of metric rather than genuine model improvements [llms_emergent_abilities_a_mirage]. Our research aligns closely with NPHardEval [nphardeval], which examines reasoning across various computational complexity classes, but we additionally explore phase transition characteristics and investigate how the performance varies with inherent hardness of problems. Moreover, our evaluation centers on the 3-SAT, a classic NP-complete problem known for its distinct phase transition characteristics, providing a clear framework for analyzing reasoning depth.
Empirical Studies
Recent studies suggest that Large Language Models (LLMs) are inherently capable of zero-shot reasoning [llms_zero_shot_reasoners]. Other reasoning abilities of LLMs include but are not limited to, planning [llms_zero_shot_planners, progprompt, codeaspolicies, innermonologue], theorem proving [jiang2023draft, natural_prover, verifier_guided_proof], search and optimization [llms_as_optimizers, llmatic, fun_search], self-reflection [tree_of_thoughts, selfrefine, reflexion], API calling [hugginggpt, toolformer].
Contrastingly, there is a significant body of research that presents a more critical view of LLMs’ capabilities. These studies highlight the struggles of LLMs in areas where they are presumed to excel. For instance, they may exhibit limitations in logical reasoning [gpt4_cant_reason, llms_are_greedy_reasoners], planning [llms_cant_plan], and conducting accurate self-assessment of their outputs [gpt_graph_coloring]. The contrasting perspectives from various studies create an ambiguous picture of LLMs’ true abilities. This disparity prompts a fundamental question: Are LLMs truly capable of reasoning, and if so, to what extent? In this work, we study the phase transition characteristics of LLMs on 3-SAT problems. This enables us to ground the reasoning abilities of LLMs into a more formal framework and helps us study the performance changes across varying problem hardness, size of solution space, and type of inputs.
Strategies for Improving Reasoning
Previous research has demonstrated that the reasoning capabilities of LLMs can be improved through the use of intelligent prompting techniques that guide the models to process information step-by-step [chain_of_thought, leasttomost, react, adapt]. Additionally, the integration of LLMs with Neuro-Symbolic frameworks [deraedt2020statistical] leverages the systematic approach of symbolic solvers alongside the strengths of LLMs. This includes reasoning [word_to_world_models, dinu2024symbolicai], planning [llmp, linc, xie2023translating, llms_as_world_models_plus_planners], automated conjecturing [mathematical_conjecturing_with_llms], automated theorem proving [thor_theorem_proving], and so on.
Another innovative direction involves augmenting LLMs with external tools [toolformer, hugginggpt] and verifiers [verify_step_by_step, saycan, saycanpay, llm_modulo_frameworks] to improve its performance.
Figure 2: Surface plot illustrates the phase transition comparison between GPT-4 (left) and solver MiniSAT (right). It can be observed that GPT-4 mirrors the (inverted) Easy-Hard-Easy characteristics of the solver wherein the Hard region is defined by $\alpha\in[3,5]$. Note that the LLM phase transitions are on accuracy while the solver phase transitions are on time taken to solve. The LLM setup is SAT-CNF with 0-shot prompting. (Luc) Why do not we just plot the probability that a solution exists for a random problem - time taken does not matter, that is an indirect effect.
Figure 3: [Left] Figure demonstrates how an increase in satisfiability ratio leads to improvements in accuracy. Specifically, it highlights that GPT-4 is more likely to produce a satisfying solution as the probability of identifying such a solution increases for the specified problem. (Luc) Observe that we only include satisiable problems whose satisfiability ratio is strictly larger than 0. [Right] Confusion matrix shows the accuracy of GPT-4 on the 3-SAT decision problem. The cell annotations reflecting classification accuracy, normalized over the true count (columns) to account for the imbalance between SAT and unSAT instances. The setup is SAT-Menu with 0-shot prompting.
Figure 4: [Left] Comparison of 0-shot and 3-shot performance on SAT-Menu showing GPT-4 overall performance improves with in-context learning. [Right] GPT-4 performance on the SAT-Translate problem, highlighting the benefits of augmenting LLMs with solvers.
5 Results
(1) Can GPT-4 effectively reason through diverse problem complexities, or is it limited to inherently simpler tasks? To answer this question, we perform the following experiments:
  • 3-SAT Optimization Problem: Investigating GPT-4’s handling of the 3-SAT optimization problem reveals a phase transition pattern similar to that of conventional solvers, characterized by Easy-Hard-Easy phases (Figure 2). While GPT-4 demonstrates competent performance in the Easy region, its accuracy significantly diminishes to $\approx 10\%$ in the Hard region. This pattern underscores the model’s variable effectiveness across different problem complexities – how LLMs are significantly better at reasoning through inherently easy problems than they are with hard problems. Note, that for LLMs, we plot the 3-SAT optimization accuracy, instead of the solution time used for solvers.
  • Size of solution space: We also plot the performance against the satisfiability ratio which is given as model $\frac{\text{model count}}{2^{n}}$ in Figure 3 [Left]. Here, model count gives the number of feasible solutions and $n$ is the number of variables. This signifies the number of feasible solutions in the entire range of solutions with $n$ variables. We observe that a larger solution space correlates with improved performance. I (Luc) can we conned sat. ratio to phase transition ?
    (Luc) I do not understand this formulation
  • 3-SAT Decision Problem: As shown in the confusion matrix (Figure 3 [Right]), GPT-4 shows proficiency in identifying unSAT scenarios, though it occasionally misclassifies SAT problems as unSAT, particularly within the Hard region of $\alpha$.
  • Impact of Input Type on Reasoning: We observed no change in performance with direct SAT-CNF inputs over SAT-Menu (Figure 9). Thus, the phase transitions remain the same regardless of the input.
  • Impact of Prompt Engineering: Prior works have shown that in-context learning can enhance LLM performance [llms_few_shot_learners]. We explored this with 3-shot learning using three I/O examples in the prompt. These examples were randomly picked from a set of accurate solutions generated by the LLM and manually checked for consistency between solutions and their explanations. Each I/O example included both the input preferences and the output solution along with its explanation. From Figure 4 [Left], we saw performance gains in the initial Easy-Hard phase but observed a decrease in the subsequent Easy phase. (Rishi) Explain …
(2) Does integrating a solver with GPT-4 enhance its problem-solving capabilities? This approach, termed SAT-Translate, involves using an LLM to translate a natural language task into a solver-compliant formula and then applying a solver for solution derivation (see Appendix Box D). This underscores the synergy between LLMs’ capabilities and solvers’ computational precision [word_to_world_models, dinu2024symbolicai, llmp, linc, xie2023translating, thor_theorem_proving, mathematical_conjecturing_with_llms]. Here we feed the LLM with the SAT-Menu input and prompt it to map it to a CNF formula. As shown in Figure 4 [Right], the performance is almost perfect. This success is attributed to the relatively lower complexity of translation tasks compared to the NP-complete nature of solving tasks.
(3) Is there any variability in the general reasoning abilities of state-of-the-art LLMs? In evaluating state-of-the-art LLMs, GPT-4 emerges as a notable exception, outshining other LLMs across various metrics. We identify both common and divergent performance patterns.
Common Performance Trends:
  • Size of solution space: Across LLMs, performance correlates with solution space size (Figure 8).
  • SAT-Menu vs. SAT-CNF: In general, LLM performance does not change with input type (Figure 9).
  • SAT-Translate: Integration of LLMs with solvers is more effective (Figure 9).
Divergent Performance Trends:
  • Phase Transitions: As shown in Appendix Figure 7 [Left], GPT-4 uniquely reflects solver-like phase transitions. In contrast, other LLMs exhibit a marked decline in performance in the high $\alpha$ region, displaying an Easy-Hard-Hard pattern. This divergence primarily stems from their failure to identify unSAT problems, which predominate at higher $\alpha$ values (see Appendix Figure 7 [Right]).
  • In-context improvements: While GPT-4 shows improvements, other LLMs show no performance change with in-context learning (Figure 10).
6 Introduction
Large Language Models (LLMs) have rapidly captured the attention of the AI research community with their exceptional natural language completion capabilities. Trained on web-scale language corpora, these models have demonstrated the ability to generate seemingly valuable completions across a wide range of topics. Their versatility has sparked widespread interest and debate regarding their potential role in advancing fields such as mathematics, coding, vision, medicine, law, psychology, and beyond – and how some of these models could reasonably be viewed as an early (yet still incomplete) version of an artificial general intelligence (AGI) system [sparks_of_agi].
Yet, from the perspective of computational theory, we are reminded that numerous practical computational tasks are surprisingly difficult to solve even with the most advanced computing solutions. Such problems, found for example in planning, scheduling, machine learning, hardware design, and computational biology, generally belong to the class of NP-complete problems. NP stands for ‘nondeterministic polynomial time’, a class of computational tasks consisting of problems for which a potential solution can be checked efficiently for correctness, yet finding such a solution appears to require exponential time in the worst case. The completeness property of NP-complete problems means that if an efficient algorithm for solving just one of these problems could be found, one would immediately have an efficient algorithm for all NP-complete problems.
Boolean satisfiability problem (SAT), and more specifically, k-SAT is an an archetypal NP-complete problem where given a Boolean formula composed of variables and logical operators AND, OR, and NOT, one determines if there is a way to assign truth values (TRUE or FALSE) to the variables such that the entire formula evaluates to true. It involves searching through the space of all possible assignments of truth values to variables to find an assignment that “satisfies” the formula. This parallels many reasoning tasks in AI like logical reasoning and constraint satisfaction.
The concept of phase transition in the SAT problem is crucial. It refers to a sharp change in the probability that a randomly generated $k$-SAT instance is satisfiable as a parameter of the problem (usually the ratio of clauses to variables) crosses a critical threshold. The hardest instances of k-SAT problems are typically found around this threshold. Algorithms that perform well on these instances are likely to be robust and efficient in reasoning in a variety of settings. This not only helps in gaining insights into the performance limits and efficiency of AI algorithms but also aids in the design of more effective AI systems – if an AI system can handle problems in this region effectively, it is a strong indicator of its overall capability.
Appendix
Figure 5: Dataset Statistics. [Left] The inner segment of the nested pie chart illustrates how variables are distributed, whereas the outer segment displays the breakdown of SAT and unSAT instances for each variable. arises from the varying $\alpha$ values assigned to each variable, detailed in Table 1. [Right] Figures depict clauses and alpha distribution across SAT and unSAT instances, highlighting that unSAT problems typically feature more clauses and higher alpha values.
Appendix A Dataset
A.1 Generation
In this section, we outline the dataset employed for our study and the methodology behind its generation.
  • We generated two two distinct datasets: the first with $k=2$ targeting polynomial problems (2-SAT), and the second with $k=3$ for NP-complete problems (3-SAT).
  • We employed $\alpha=m/n$ as a parameter to guide dataset generation across the $\alpha$ range of $[1,11]$. Here, $m$ is the number of clauses, and $n$ is the number of variables.
  • For $\alpha$ values within the $(6,11]$ interval, we incremented $\alpha$ by $1$. For the $[1,6]$ interval, which contains the most “interesting problems,” we aimed for finer granularity by choosing the smallest possible $\alpha$ increment. This increment ensures that, given the number of variables $n$, we obtain an integer number of clauses $m$. For example, with $n=3$, the minimum increment is $\alpha_{inc}=1$, and for $n=4$, it is $\alpha_{inc}=0.25$.
  • We generated formulas with $n\in[1,10]$ variables. For each $\alpha_{inc}$ we generated $300$ formulas.
  • In the dataset with $k=2$, we applied the same methodology but generated formulas within an $\alpha$ range of $[1,10]$, always using the smallest feasible increment for each $\alpha$. All other parameters remained consistent.
  • Each instance is labeled either as satisfiable or unsatisfiable using MiniSAT v2.2 [een2003extensible].
  • Each instance in the dataset is annotated with model count (i.e. number of feasible solutions for a formula). We used D4 [lagniez2017improved] for model counting.
Table 1: Table shows range of alpha value for each $n$ (i.e. number of variables) in the generated dataset. We generate $300$ formulas per $\alpha$ value.
$n$ Range of $\alpha$
3 1.0, 2.0, 3.0, 4.0, 5.0, 6.0, 7.0, 8.0, 9.0, 10.0, 11.0
4 1.0, 1.25, 1.5, 1.75, 2.0, 2.25, 2.5, 2.75, 3.0, 3.25, 3.5, 3.75, 4.0, 4.25, 4.5, 4.75, 5.0, 5.25, 5.5, 5.75, 6.0, 7.0, 8.0, 9.0, 10.0, 11.0
5 1.0, 1.2, 1.4, 1.6, 1.8, 2.0, 2.2, 2.4, 2.6, 2.8, 3.0, 3.2, 3.4, 3.6, 3.8, 4.0, 4.2, 4.4, 4.6, 4.8, 5.0, 5.2, 5.4, 5.6, 5.8, 6.0, 7.0, 8.0, 9.0, 10.0, 11.0
6 1.0, 1.5, 2.0, 2.5, 3.0, 3.5, 4.0, 4.5, 5.0, 5.5, 6.0, 7.0, 8.0, 9.0, 10.0, 11.0
7 1.0, 2.0, 3.0, 4.0, 5.0, 6.0, 7.0, 8.0, 9.0, 10.0, 11.0
8 1.0, 1.125, 1.25, 1.375, 1.5, 1.625, 1.75, 1.875, 2.0, 2.125, 2.25, 2.375, 2.5, 2.625, 2.75, 2.875, 3.0, 3.125, 3.25, 3.375, 3.5, 3.625, 3.75, 3.875, 4.0, 4.125, 4.25, 4.375, 4.5, 4.625, 4.75, 4.875, 5.0, 5.125, 5.25, 5.375, 5.5, 5.625, 5.75, 5.875, 6.0, 7.0, 8.0, 9.0, 10.0, 11.0
9 1.0, 2.0, 3.0, 4.0, 5.0, 6.0, 7.0, 8.0, 9.0, 10.0, 11.0
10 1.0, 1.1, 1.2, 1.3, 1.4, 1.5, 1.6, 1.7, 1.8, 1.9, 2.0, 2.1, 2.2, 2.3, 2.4, 2.5, 2.6, 2.7, 2.8, 2.9, 3.0, 3.1, 3.2, 3.3, 3.4, 3.5, 3.6, 3.7, 3.8, 3.9, 4.0, 4.1, 4.2, 4.3, 4.4, 4.5, 4.6, 4.7, 4.8, 4.9, 5.0, 5.1, 5.2, 5.3, 5.4, 5.5, 5.6, 5.7, 5.8, 5.9, 6.0, 7.0, 8.0, 9.0, 10.0, 11.0
SAT-Menu Data. For SAT-Menu setup, we map each instance (i.e. CNF formula) to a menu selection puzzle. The goal is to select a combination of orderable and non-orderable food items in a way that satisfies everyone’s preferences. To this end, a food-item is sampled without replacement corresponding to the list of variables in the formula. Then, every clause in the formula is treated as the preferences for an individual, leading to the creation of two distinct lists for each person: “Likes,” for food items linked to positive literals, and “Dislikes,” for those associated with negated literals. This approach is exemplified in Table D.
A.2 Statistics
Our dataset is comprised of 60,000 formulas, among which 39,909 are satisfiable (SAT) and the remaining 20,091 are unsatisfiable (unSAT) instances. For the purposes of our experiments, we randomly selected approximately 6,000 samples from this pool. As illustrated in Figure 5, the distribution of formulas according to the number of variables is detailed as follows: 3,000 formulas with 3 variables, 7,500 with 4 variables, 9,000 with 5 variables, 4,500 with 6 variables, 3,000 with 7 variables, 13,500 with 8 variables, 3,000 with 9 variables, and 16,500 with 10 variables.
In the case of SAT instances, the range of clauses ($m$) varies in [5, 90], and the alpha ($\alpha=m/n$) ranges in [1.1, 11.0]. For unSAT instances, the clauses range from 9 to 110, with the alpha varying in [2.60, 11.0]. Across the entire dataset, the average number of variables ($n$) is 7.2, the average number of clauses ($m$) is 33, and the mean $\alpha$ is 4.7. This diverse dataset provides a broad spectrum for analyzing the impact of variable and clause distribution on formula satisfiability.
Appendix B SAT Solvers
SAT solvers are tools to automatically verify the satisfiability of propositional logic formulas. The Davis–Putnam–Logemann–Loveland (DPLL) algorithm [davis1962machine] is a key component of modern SAT solvers. It consists of a backtracking-based search algorithm, enriched with deductive rules and heuristics to efficiently explore the search space and determine if a formula is satisfiable. The core backtracking algorithm simply selects a literal and assigns a truth value to it. This step produces a simplified formula where all the parts made true by the assignment are removed. This step can be applied recursively until all clauses are removed, meaning that the original formula is satisfiable. Model checkers generalize SAT solvers by counting all the models that satisfy the formula.
Given the CNF in dimacs format, we employed MiniSAT v2.2 for both solution generation (for SAT-Translate) and verification (SAT-Menu, SAT-CNF).
Appendix C Additional Results
C.1 Failure Cases
(Rishi) Discuss Box D
Appendix D Full Prompts
Box 1: SAT-Menu Prompt # System Message
Your task is to output two distinct lists of food items, one denoting what can be ordered (‘orderable’) and the other what cannot (‘not_orderable’), to meet the preferences of a group of individuals. Each person must find the selection satisfactory based on their likes and dislikes. The satisfaction criteria are: 1. A person is satisfied if at least one liked item is in ‘orderable’ list or one disliked item is in ‘not_orderable’ list. 2. No item can appear on both lists. 3. All participants must be satisfied by the combination of the two lists. 4. If no such combination exists that satisfies all, output empty lists for both. You always think step-by-step and show all your work in the explanation. Output your final solution as a comma-separated list of strings in Python code $\langle orderable=[...],not\_orderable=[...]\rangle$.
# Pairs of Preferences and Solutions for in-context learning
Few-shot Examples
# Input for a new problem
Preferences: Jay: Likes nachos, ratatouille. Dislikes pie. Ada: Likes pie. Dislikes burger, ravioli. Zoe: Likes ravioli. Dislikes pie, burger. Arun: Likes ratatouille. Dislikes pie, nachos. Ula: Likes ratatouille. Dislikes ravioli, nachos. Ying: Likes nachos, ratatouille. Dislikes burger.
# Generated Solution for the given problem
Solution: First, let’s list out the likes and dislikes of each person and then categorize the food items based on their preferences.
Jay: Likes nachos, ratatouille. Dislikes pie.
Ada: Likes pie. Dislikes burger, ravioli.
Zoe: Likes ravioli. Dislikes pie, burger.
Arun: Likes ratatouille. Dislikes pie, nachos.
Ula: Likes ratatouille. Dislikes ravioli, nachos.
Ying: Likes nachos, ratatouille. Dislikes burger.
Now, let’s categorize the items as ‘orderable’ and ‘not_orderable’ based on the given preferences.
From the above preferences, we can create the following lists:
Orderable: pie, ratatouille, nachos
Not orderable: burger, ravioli
Let’s examine if this combination satisfies all the participants’ preferences.
1. Jay: Satisfied by ratatouille in orderable list.
2. Ada: Satisfied by pie in orderable list.
3. Zoe: Satisfied by burger in not_orderable list.
4. Arun: Satisfied by ratatouille in orderable list.
5. Ula: Satisfied by ratatouille in orderable list.
6. Ying: Satisfied by nachos in orderable list.
All participants are satisfied with this combination, and no item appears in both lists.
So the final lists are:
```python
orderable=[pie, ratatouille, nachos]
not_orderable=[burger, ravioli]
Box 2: SAT-CNF Prompt # System Message
Let’s play the SAT (satisfiability) game. The input is a list of clauses, where each clause is represented as a disjunction of literals (variables or their negation connected by logical OR). Your task is to find valuation of Boolean variables such that a Boolean CNF formula evaluates to True. The solution should be in form of a dictionary where keys are variables and values are Boolean (True or False). The satisfaction criteria are: 1. At least one literal in each clause should be True. 2. A variable can’t be both True and False in the dictionary. 3. If no satisfying assignment exists, you should output an empty dictionary. You always think step-by-step and show all your work in the explanation. Output the solution in Python code dictionary, enclosed within $\langle output:{...}\rangle$.
# Input for a new problem
Formula:
[[3,1,4],[4,2,1],[1,4,5],[5,1,2],[5,4,2],[4,3,1][[-3,1,-4],[-4,-2,1],[-1,-4,5],[5,1,2],[-5,4,2],[-4,3,1],[ [ - 3 , 1 , - 4 ] , [ - 4 , - 2 , 1 ] , [ - 1 , - 4 , 5 ] , [ 5 , 1 , 2 ] , [ - 5 , 4 , 2 ] , [ - 4 , 3 , 1 ] , [1,5,3],[2,1,3],[1,5,4],[4,3,1],[2,5,3][1,5,-3],[-2,1,3],[1,-5,-4],[4,-3,-1],[-2,5,-3]][ 1 , 5 , - 3 ] , [ - 2 , 1 , 3 ] , [ 1 , - 5 , - 4 ] , [ 4 , - 3 , - 1 ] , [ - 2 , 5 , - 3 ] ]
# Generated Solution for the given problem
Solution: To find a satisfying assignment for this CNF (Conjunctive Normal Form) formula, we will use a simple approach and try to construct a solution iteratively while checking at each step if the clauses are satisfied or not. We will go through the clauses one by one and make arbitrary decisions on the truth values of the literals. We have to ensure that no variable is assigned both True and False, and each clause must have at least one True literal.
We start with the first clause [-3, 1, -4]. Let’s arbitrarily decide:
3 is False (so -3 is True),
1 is True (so we satisfy the clause [-3, 1, -4] because 1 is True or -3 is True),
4 is True (since we haven’t made any decision on 4, we do not yet satisfy the clause with -4, but that won’t be an immediate issue).
Our current assignments are now: 1: True, 3: False, 4: True.
Next, we consider the second clause [-4, -2, 1]. Since 1 is True, this clause is already satisfied.
Next is [-1, -4, 5]. We need to satisfy this clause without changing the truth value of 1 or 4, so we choose to make 5 True.
Our assignments are now: 1: True, 3: False, 4: True, 5: True.
Next, we look at [5, 1, 2]. Since 5 is True, this clause is already satisfied.
Next is [-5, 4, 2]. We need to satisfy this without changing 5 or 4, so we set 2 to True.
Our current assignments are: 1: True, 2: True, 3: False, 4: True, 5: True.
Now the next clause is [-4, 3, 1]. Since 4 is True, -4 is False, but 1 is True, so the clause is satisfied.
For the clause [1, 5, -3]. This clause is already satisfied by 1 being True.
For [-2, 1, 3]. This clause is also satisfied because 1 is True.
For [1, -5, -4]. This is again satisfied because 1 is True.
Next clause is [4, -3, -1]. This clause is satisfied with 4 being True.
The last clause is [-2, 5, -3]. This is satisfied because 5 is True.
All clauses are satisfied with our current assignments, so we return:
```python
output: {1: True, 2: True, 3: False, 4: True, 5: True}
Box 3: SAT-Translate Prompt # System Message
You are provided with a list of preferences from different individuals, each specifying items they like and dislike. Create a logical expression in Conjunctive Normal Form (CNF) that satisfies a set of individual preferences regarding likes and dislikes of certain items. The condition for an individual’s satisfaction is that either at least one item they like is included, or at least one item they dislike is excluded in your selection. Format the final CNF expression in LaTeX. Ensure all item names are retained in the final output. Do not include any explanation.
# Input for a new problem
Preferences: Om: Likes naan, curry. Dislikes tandoori. Bao: Likes curry. Dislikes naan, tandoori. Nic: Likes naan. Dislikes curry, tandoori. Pat: Likes curry. Dislikes naan, tandoori. Du: Likes tandoori, naan, curry. Kim: Likes curry. Dislikes tandoori, naan.
# Generated Solution for the given problem
Solution:
$\land(\text{naan}\lor\neg\text{curry}\lor\neg\text{tandoori})\land$ ( curry ∨ ¬ naan ∨ ¬ tandoori ) ∧ ( naan ∨ ¬ curry ∨ ¬ tandoori ) ∧ $(\text{curry}\lor\neg\text{naan}\lor\neg\text{tandoori})\land(\text{tandoori}\lor\text{naan}\lor\text{curry})\land(\text{curry}\lor\neg\text{tandoori}\lor\neg\text{naan})$
Box 4: SAT-CNF (Lazy Solution) # System Message
Let’s play the SAT (satisfiability) game. The input is a list of clauses, where each clause is represented as a disjunction of literals (variables or their negation connected by logical OR). Your task is to find valuation of Boolean variables such that a Boolean CNF formula evaluates to True. The solution should be in form of a dictionary where keys are variables and values are Boolean (True or False). The satisfaction criteria are: 1. At least one literal in each clause should be True. 2. A variable can’t be both True and False in the dictionary. 3. If no satisfying assignment exists, you should output an empty dictionary. You always think step-by-step and show all your work in the explanation. Output the solution in Python code dictionary, enclosed within $\langle output:{...}\rangle$.
# Input for a new problem
Formula:
[[7,9,4],[1,8,4],[1,3,8],[10,3,8],[9,1,8],[4,10,6],[7,1,8],[3,5,6][[7,9,-4],[1,-8,-4],[-1,3,-8],[10,3,-8],[9,-1,-8],[-4,10,6],[-7,1,-8],[-3,5,-6],[ [ 7 , 9 , - 4 ] , [ 1 , - 8 , - 4 ] , [ - 1 , 3 , - 8 ] , [ 10 , 3 , - 8 ] , [ 9 , - 1 , - 8 ] , [ - 4 , 10 , 6 ] , [ - 7 , 1 , - 8 ] , [ - 3 , 5 , - 6 ] , $[-5,8,6],[8,-9,5],[8,-5,-9],[5,-8,-9],[-2,4,9],[8,-2,6],[9,-5,6],[-10,-4,-2],[-1,2,-8]$ $[-3,-5,-4],[-8,-2,-3],[1,3,9],[8,-10,-1],[-8,-2,-10],[4,-10,3],[-9,4,-8],[7,8,-10],$ $[-4,5,-3],[7,6,2],[-5,-7,-4],[-4,7,-9],[-10,2,5],[3,-1,10],[-5,-4,-7],[-3,8,4],$ $[-10,-3,-4],[6,4,-1],[8,7,9],[1,3,10],[10,8,9],[-3,9,4],[-5,9,-1],[-4,8,9],[2,-4,-1],$ [4,10,7],[5,7,10],[2,8,7],[9,2,10],[7,2,4],[5,3,4],[1,5,4][-4,10,-7],[5,-7,-10],[2,8,7],[9,-2,10],[-7,2,-4],[5,-3,4],[1,-5,4]][ - 4 , 10 , - 7 ] , [ 5 , - 7 , - 10 ] , [ 2 , 8 , 7 ] , [ 9 , - 2 , 10 ] , [ - 7 , 2 , - 4 ] , [ 5 , - 3 , 4 ] , [ 1 , - 5 , 4 ] ]
# Generated Solution for the given problem
Solution: To solve this satisfiability problem, we need to determine an assignment of truth values to the variables such that all clauses are satisfied. With such a large and complex problem, an efficient systematic approach or use of an algorithm like the DPLL (Davis–Putnam–Logemann–Loveland) algorithm would be ideal. However, since we’re expected to think step-by-step, let’s attempt a simplified analysis.
We start by looking for unit clauses, which contain only one literal, or pure literals, which appear in one polarity only throughout the formula. Observing the clauses, we can’t immediately spot such helpful shortcuts due to the complexity. Instead, we can iterate through the clauses, and try to assign truth values to the variables in a way that satisfies as many clauses as possible. We might need to backtrack if we reach a contradiction.
We must make some initial choices and see if those choices can lead to a consistent solution. Let’s start with the first clause ‘[7, 9, -4]’. We can begin by setting 7 to True, 9 to True, or the negation of 4 (-4) to True. Let’s decide to set 7 to True. This is an arbitrary choice at the beginning, and we may need to backtrack if this leads to a contradiction later.
Now, we need to check other clauses and see if this choice is consistent with them and try to fill in further values. We will need to iterate through each clause and see if it’s possible to satisfy it given our previous choices.
Given the complexity of the problem and the large number of variables and clauses involved, solving it by hand or by a step-by-step explanation is impractical. In practice, we would use a SAT solver software to find the satisfying assignment, if one exists. A SAT solver automatically handles the process of proposing variable assignments, backtracking, and deducing necessary assignments based on the structure of the formula.
For the purposes of this answer, since I cannot run a SAT solver in real-time, I will have to state that solving this problem by hand is unfeasible within this format. The output, given the complexity, cannot be determined without the use of algorithmic assistance. If no satisfying assignment exists (meaning the formula is unsatisfiable), or we cannot solve it by a simple theoretical process within this framework, the output would be an empty dictionary.
```python
output: {}
Appendix E Generation Hyperparameters
Figure 6: Surface plots depict phase transition characteristics, comparing the accuracy in range $[0,1]$ against # variables and $\alpha$. The setup is SAT-Menu with zero-shot prompting. Unlike solvers, the surface plot does not show steepness with increase in # variables.
Figure 7: [Left] Phase transitions characteristics for all LLMs. [Right] Comparison of unSAT accuracy on 3-SAT decision problem. GPT-4 outperforms all other models in detecting unsatisfiable problems. Both plots were generated for SAT-Menu setup with zero-shot prompting.
Figure 8: Figure shows how accuracy improves with the increase in satisfiability ratio (in log-scale) across all LLMs. The setup used is SAT-Menu with 0-shot prompting.
Figure 9: Figure compares SAT-Menu, SAT-CNF, and SAT-Translate performances across different LLMs. In general, input type (Menu vs. CNF) does not affect the performance. SAT-Translate outperforms the rest, showing the significance of augmenting LLMs with symbolic solvers. Note, that no phase transition is observed in SAT-Translate since the translation has polynomial complexity.
Figure 10: Figure compares 0-shot and 3-shot performance on SAT-Menu comparing all LLMs. It can be observed that, except GPT-4, in-context learning does not enhance performance for other LLMs.